kennedywoods:

What is the value of x? 34x+5−12x=3 −8 −12 12 8

kennedywoods:

please help!!!!

AnimeLuver10023:

22x + 5 = 3 22x = 3 - 5 22x = -2 x = -2/22 x = -1/11

Vocaloid:

@kennedywoods is there a typo in your original problem?

Vocaloid:

@kennedywoods Closed due to answered.